### v2

@@ Abstract
1 Introduction
@@
Temporal logics are widely used as specification languages for formal program verification and synthesis. They are also commonly used by the AI community for reasoning about actions and planning, such as expressing temporal constraints in task planning. One such logic, Linear Temporal Logic (LTL)[22], has been widely used in many contexts. It extends standard propositional logic with additional operators to express temporal constraints. It is simple to understand and easy to use effectively. Moreover, it has an attractive balance between formal expressiveness, i.e., the types of properties that can be specified by a formula, and the complexity of performing basic operations on formulas, such as satisfiability checking, which is a standard measure for determining the computational tractability of a formal logic.
@@
A more recently developed temporal logic is Linear Dynamic Logic (LDL)[14]. LDL offers increased formal expressiveness, which makes it useful in cases when LTL’s lack of expressiveness is an obstacle, which is not uncommon in practice [14]. LDL maintains a comparable ease of use to LTL. It combines propositional logic with the use of path expressions, which are regular expressions used to express temporal constraints. Since most computer scientists are already familiar with regular expressions, it is easy for them to learn to use path expressions effectively.
@@
The semantics of temporal logics may be defined over infinite traces or finite traces. The former allows for reasoning over an unbounded length of time, while the latter allows for reasoning over a finite time horizon. LTL was originally defined to operate over infinite traces[22], but more recent variants of LTL have been defined over finite traces and have proven to be useful in practice[1, 21, 14]. Given the usefulness of finite-trace temporal logics and the increasing degree of interest in them, we limit our scope here to considering finite-trace variants of temporal logics and leave infinite-trace variants to future work.
@@
A limitation of LDLf is that its only published NFA construction, which is provided in [14] and used to check satisfiability, is not well defined for all syntactically valid LDLf formulas. There is a class of path expressions that if present in an LDLf formula will cause the automaton construction to fail. In particular, if a path expression contains a Kleene star operator whose operand is an expression that can be of length 0, then following the construction in [14] results in infinite recursion while building the transition function.
@@
The main contribution of this work is the introduction of Automata Linear Dynamic Logic on Finite Traces (ALDLf), a variant of LDLf, and showing that satisfiability for ALDLf is in PSPACE. ALDLf uses path automata, which are nondeterministic finite automata (NFA) used to express temporal constraints. Using path automata rather than path expressions allows for an automaton construction that is valid for all ALDLf formulas. Moreover, by converting path expressions to path automata using classic algorithms such as Thompson’s construction for converting regular expressions to NFA [27], the automaton construction for ALDLf can also be used for all LDLf formulas. Additionally, ALDLf provides for the direct expression of past constraints, in contrast to LDLf, which can only make direct claims about the present and future. As part of the means of establishing ALDLf satisfiability, we also introduce a novel variant of the two-way alternating automaton on finite words (2AFW) that, in addition to conventional finite-run acceptance conditions, uses a Büchi-like acceptance condition on infinite runs to allow some infinite runs to be accepting.
@@
ALDLf is equivalent in expressiveness to LDLf, which is equivalent to Monadic Second-Order Logic, while LTL is equivalent to Monadic First-Order Logic. Since LTL satisfiability is PSPACE-complete and this work demonstrates that ALDLf satisfiability is in PSPACE, ALDLf provides greater expressiveness and succinctness than LTLf without increasing the complexity of satisfiability. The motivation to add automata connectives to LDLf is two fold. First, modern industrial property specification languages such as SystemVerilog Assertions (SVA) allowed assertions to contain local variables, allowing the expression of local state, which is not allowed in LDLf. Second, on the theoretical side, as finite-state automata are exponentially more succinct that regular expressions [16], our investigation here explores the succinctness/complexity terrain for finite-horizon temporal logics, analogously to [30].
1.1 Outline
@@ 2 Preliminaries
LTLf, which is a variant of LTL defined for finite traces, is introduced in [14]. Its formulas are generated from the same grammar as standard LTL, which extends propositional logic with two additional operators, $X$ (“next”) and $\mathcal{U}$ (“until”):
$\varphi::=P\mid(\neg\varphi)\mid(\varphi_{1}\wedge\varphi_{2})\mid(X\varphi)\mid(\varphi_{1}\mathcal{U}\varphi_{2})$
The definition semantics is defined as follows. Let $\pi$ be a trace of finite length and $|\pi|$ be the length of $\pi$. The satisfaction of an LTLf formula $\varphi$ at time point $i$ on $\pi$, written as $\pi,i\models\varphi$, is inductively defined as follows:
  1. 1.
@@
We use the abbreviations $\varphi_{1}\vee\varphi_{2}\equiv\neg(\neg\varphi_{1}\wedge\neg\varphi_{2})$, $\varphi_{1}\rightarrow\varphi_{2}\equiv\neg\varphi\vee\varphi_{2}$, and $\varphi_{1}\leftrightarrow\varphi_{2}\equiv(\varphi_{1}\rightarrow\varphi_{2})\wedge(\varphi_{2}\rightarrow\varphi_{1})$. If $\varphi$ is a formula then $AP(\varphi)$ denotes the set of atomic propositions that occur in $\varphi$. If $\varphi$ is a formula then $\operatorname{nnf}(\varphi)$ an equivalent formula in negation normal form. We use $|\pi|$ to denote the length of $\pi$, $\pi_{i}$ to denote the $i$th element of $\pi$, and $[\pi]$ to denote the set $\{0,\ldots,|\pi|-1\}$ of positions in $\pi$. The symbol true denotes a propositional formula that always evaluates to true and the symbol false denotes a propositional formula that always evaluates to false.
3 Related Work
@@
The use of temporal logic for use in formal program verification was first proposed by Pnueli in 1977 [22], using an adaptation of Tense Logic, which was first developed by Prior in 1957 to facilitate precise reasoning about time for use in solving philosophical problems [23]. Pnueli’s adaptation, which came to be known as Linear Temporal Logic (LTL), was tailored to have the necessary expressiveness to specify temporal dependencies in programs. Pnueli’s LTL was originally used over infinite traces. More recently LTL interpreted over finite traces (LTLf) was proposed in [14] and shown to be a useful specification language[1, 5, 11, 21, 33].
@@
Regular expressions over propositional formulas (REf), where a propositional formula $\varphi$ serves as a representation of the set of propositional interpretations that satisfy $\varphi$, are a more expressive alternative to LTLf, with regular expressions being equivalent in expressiveness to monadic second-order logic (MSO) and LTLf being equivalent in expressiveness to monadic first-order logic (FOL). Nevertheless, REf lacks direct constructs for negation and conjunction. These can be added by allowing complementation and intersection of regular expressions, but this results in nonemptiness checking being of nonelementary complexity even for star-free regular expressions.
@@
Linear Dynamic Logic on Finite Traces (LDLf) is a temporal logic that extends propositional logic with operators that use regular expressions to express temporal constraints [14]. It has the formal expressiveness of REf while having constructs for negation and conjunction. Notably, satisfiability checking for LDLf formulas is PSPACE-complete, the same as for LTLf. Thus, LDLf is more expressive than LTLf at no additional cost[14].
4 Introducing ALDLf
@@
Automata Linear Dynamic Logic on Finite Traces (ALDLf) is a variant of LDLf that extends propositional logic with operators that use nondeterministic finite automata (NFA) to express temporal constraints. It also provides direct constructs for past modalities. The use of NFAs rather than regular expressions offers two advantages. The first advantage is that NFAs are exponentially more succinct than regular expressions[15, 16]. Satisfiability checking of ALDLf formulas is PSPACE-complete, so this succinctness comes at no cost in complexity. The second advantage is that sometimes NFAs can be a more convenient form of expression than regular expressions. It is sometimes easier in practice to describe behavior using a state-based formalism, as can be observed by the the use of such formalism in industrial language. For example, SystemVerilog Assertions (SVA) uses local variables[32], which are a form of state.
@@
Moreover, ALDLf and its satisfiability-checking procedure can be used for LDLf, by converting an LDLf formula to an equivalent ALDLf formula. This can be done by converting each occurrence of regular expression in the LDLf formula to an equivalent NFA using a standard algorithm, e.g., Thompson’s construction[27]. Having an alternative method to perform LDLf satisfiability checking to the one provided in [14] is advantageous because there are syntactially valid LDLf formulas for which the construction of the transition function for the AFW provided in [14] may not terminate. A simple example is $\varphi=\langle p^{**}\rangle false$, where $p$ is an atomic proposition. It is easy to see that following the AFW transition function construction provided in Theorem 15 of [14] on $\varphi$ results in a loop. Nevertheless, converting $\varphi$ to an ALDLf formula and performing the satisfiability checking procedure provided here is well founded.
@@ 4.2 ALDLf Semantics
If the conditions in 4.1 below hold, then an ALDLf formula is true within a trace $\pi\in(2^{AP})^{+}$ at position $i$, where $0\leq i<|\pi|$, (denoted $\pi,i\models\varphi$), and an ALDLf path automaton $(R,T,\Delta,r,G)$ is satisfied by the subsequence of $\pi$ from position $i$ to position $j$ (denoted $\pi,i,j\models(R,T,\Delta,r,G)$ if the conditions in 4.2 holds:
@@
Intuitively, a path automaton divides a trace into a prefix and a remainder (there may be more than one such division). In the case of a formula of the form $\langle U\rangle\varphi$, a trace $\pi$ will satisfy this formula just in case there exists a prefix of $\pi$ that is in the language of $U$ such that the remainder of $\pi$ is in the language of $\varphi$. In the case of a formula of the form $[U]\varphi$, a trace $\pi$ will satisfy this formula just in case for every prefix of $\pi$ that is in the language of $U$ the remainder of $\pi$ is in the language of $\varphi$. Note that $\langle U\rangle\varphi\equiv\neg[U]\neg\varphi$ and $[U]\varphi\equiv\neg\langle U\rangle\neg\varphi$.
A trace $\pi$ is in the language of an ALDLf formula $\varphi$, denoted $\pi\models\varphi$, if $\pi,0\models\varphi$. If $\zeta$ is a propositional ALDLf formula, the notation $\pi_{i}\models\zeta$ may be used as a synonym for $\pi,i\models\zeta$, in order to emphasize the fact that the truth value of $\zeta$ at $i$ depends only on the contents of $\pi_{i}$. An ALDLf formula $\varphi$ is in negation normal form if all instances of the negation operator ($\neg$) that occur in $\varphi$, including those that occur in tests that are included in a path automaton’s alphabet, are applied to atomic propositions. An ALDLf formula can be converted to negation normal form by replacing all instances of abbreviations with their definitions, converting all instances of $\neg[U]\varphi$ to $\langle U\rangle\neg\varphi$ and all instances of $\neg\langle U\rangle\varphi$ to $[U]\neg\varphi$, applying DeMorgan’s rules to make all negations appear as inwardly nested as possible, and then eliminating all double negations.
4.3 Fischer-Ladner Closure for ALDLf
As we have just seen, the satisfaction relation for ALDLf is defined recursively, i.e., the satisfaction of a formula is defined in terms of the satisfaction of other formulas with the exception of some base cases. This is similar to how satisfaction is defined for LTL. Our strategy to translate ALDLf formulas to automata will be to use states of the automaton to represent formulas that are ”reachable” during the recursive evaluation of satisfaction, which is similar to how a similar translation is done for LTL. It is necessary to know in advance which formulas may need to be evaluated in order to construct the automaton. In the case of an LTL formula $\varphi$, the subformulas of $\varphi$ are sufficient. But in the case of modal logics such as ALDLf, the concept of subformula is too weak and so we need to use something more broad. The Fischer-Ladner closure, introduced in [10] for use in Propositional Dynamic Logic (PDL), is a generalization of subformulas for modal logics. Here we define the Fischer-Ladner closure for ALDLf.
@@ 4.4 Example
We can use these path automata to create a formula that the defines the language of traces in which $a$ being true anywhere in the trace implies that $a$ is true at the first position in the trace: $\varphi=[\alpha_{3}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false}$. For the purpose of explanation, we break down $\varphi$ into these subformulas:
@@
5 Two-way Alternating Automata on Finite Words (2AFW)
In order to serve as an intermediate step in the translation of formulas of ALDLf to equivalent NFAs, we use a two-way alternating automaton on finite words (2AFW). A useful history and taxonomy of two-way alternating automata is provided by Kapoutsis and Zakzok[18]. The 2AFW we define here is similar to that of Geffert and Okhotin[12]. The translation of ALDLf formula to NFA will first convert an ALDLf formula to an equivalent 2AFW and then convert the result to an equivalent NFA.

### v4

@@ Abstract.
Based on MS Thesis of Kevin W. Smith. Work supported in part by NSF grants IIS-1527668, CCF-1704883, IIS-1830549, DoD MURI grant N00014-20-1-2787, and an award from the Maryland Procurement Office.
1. Introduction
@@
Temporal logics are widely used as specification languages for formal program verification and synthesis. They are also commonly used by the AI community for reasoning about actions and planning, such as expressing temporal constraints in task planning. One such logic, Linear Temporal Logic (LTL) [Pnu77], has been widely used in many contexts. It extends standard propositional logic with additional operators to express temporal constraints. It is simple to understand and easy to use effectively. Moreover, it has an attractive balance between formal expressiveness, i.e., the types of properties that can be specified by a formula, and the complexity of performing basic operations on formulas, such as satisfiability checking, which is a standard measure for determining the computational tractability of a formal logic.
@@
A more recently developed temporal logic is Linear Dynamic Logic (LDL) [DGV+13]. LDL offers increased formal expressiveness, which makes it useful in cases when LTL’s lack of expressiveness is an obstacle, which is not uncommon in practice [DGV+13]. LDL maintains a comparable ease of use to LTL. It combines propositional logic with the use of path expressions, which are regular expressions used to express temporal constraints. Since most computer scientists are already familiar with regular expressions, it is easy for them to learn to use path expressions effectively.
@@
The semantics of temporal logics may be defined over infinite traces or finite traces. The former allows for reasoning over an unbounded length of time, while the latter allows for reasoning over a finite time horizon. LTL was originally defined to operate over infinite traces [Pnu77], but more recent variants of LTL have been defined over finite traces and have proven to be useful in practice [vDAPS09, PBvdA10, DGV+13]. Given the usefulness of finite-trace temporal logics and the increasing degree of interest in them, we limit our scope here to considering finite-trace variants of temporal logics and leave infinite-trace variants to future work.
@@
A limitation of LDLf is that there exist some LDLf formulas for which there is currently no published well defined NFA construction. The only published LDLf construction, which is provided in [DGV+13] and used to check satisfiability, is not well defined for these formulas. There is a class of path expressions that if present in an LDLf formula will cause the automaton construction to fail. In particular, if a path expression contains a Kleene star operator whose operand is an expression that can be of length 0, then following the construction in [DGV+13] results in infinite recursion while building the transition function.
@@
The main contribution of this work is the introduction of Automata Linear Dynamic Logic on Finite Traces (ALDLf), a variant of LDLf, and showing that satisfiability for ALDLf is in PSPACE. ALDLf uses path automata, which are nondeterministic finite automata (NFA) used to express temporal constraints. Using path automata rather than path expressions allows for an automaton construction that is valid for all ALDLf formulas. Moreover, by converting path expressions to path automata using classic algorithms such as Thompson’s construction for converting regular expressions to NFA [Tho68], the automaton construction for ALDLf can also be used for all LDLf formulas. Additionally, ALDLf provides for the direct expression of past constraints, which can be combined with present and future constraints within a single ALDLf formula. In comparison, LDLf can only directly express present and future constraints, while pure-past LDLf (PLDLf) can only directly express present and past constraints [DGDSF+20].
Additionally, ALDLf provides for the direct expression of past constraints, in contrast to LDLf, which can only make direct claims about the present and future. As part of the means of establishing ALDLf satisfiability, we also introduce a novel variant of the two-way alternating automaton on finite words (2AFW) that, in addition to conventional finite-run acceptance conditions, uses a Büchi-like acceptance condition on infinite runs to allow some infinite runs to be accepting.
@@
ALDLf is equivalent in expressiveness to LDLf, which is equivalent to Monadic Second-Order Logic, while LTL is equivalent to Monadic First-Order Logic. Since LTL satisfiability is PSPACE-complete and this work demonstrates that ALDLf satisfiability is in PSPACE, ALDLf provides greater expressiveness and succinctness than LTLf without increasing the complexity of satisfiability. The motivation to add automata connectives to LDLf is two fold. First, modern industrial property specification languages such as SystemVerilog Assertions (SVA) allowed assertions to contain local variables, allowing the expression of local state, which is not allowed in LDLf. Second, on the theoretical side, as finite-state automata are exponentially more succinct that regular expressions [GH15], our investigation here explores the succinctness/complexity terrain for finite-horizon temporal logics, analogously to [VW84].
1.1. Outline
@@ 2. Preliminaries
LTLf, which is a variant of LTL defined for finite traces, is introduced in [DGV+13]. Its formulas are generated from the same grammar as standard LTL, which extends propositional logic with two additional operators, $X$ (“next”) and $\mathcal{U}$ (“until”):
$\varphi::=P\mid(\neg\varphi)\mid(\varphi\wedge\varphi)\mid(X\varphi)\mid(\varphi$ $\mathcal{U}$ $\varphi)$
The semantics of LTLf are defined as follows. Let $\pi$ be a trace of finite length and $|\pi|$ be the length of $\pi$. The satisfaction of an LTLf formula $\varphi$ at time point $i$ on $\pi$, written as $\pi,i\models\varphi$, is inductively defined as follows:
  1. (1)
@@
We use the abbreviations $\varphi_{1}\vee\varphi_{2}\equiv\neg(\neg\varphi_{1}\wedge\neg\varphi_{2})$, $\varphi_{1}\rightarrow\varphi_{2}\equiv\neg\varphi_{1}\vee\varphi_{2}$, and $\varphi_{1}\leftrightarrow\varphi_{2}\equiv(\varphi_{1}\rightarrow\varphi_{2})\wedge(\varphi_{2}\rightarrow\varphi_{1})$. If $\varphi$ is a formula then $AP(\varphi)$ denotes the set of atomic propositions that occur in $\varphi$. If $\varphi$ is a formula then $\operatorname{nnf}(\varphi)$ an equivalent formula in negation normal form. We use $|\pi|$ to denote the length of $\pi$, $\pi_{i}$ to denote the $i$th element of $\pi$, and $[\pi]$ to denote the set $\{0,\ldots,|\pi|-1\}$ of positions in $\pi$. The symbol true denotes a propositional formula that always evaluates to true and the symbol false denotes a propositional formula that always evaluates to false.
3. Related Work
@@
The use of temporal logic for use in formal program verification was first proposed by Pnueli in 1977 [Pnu77], using an adaptation of Tense Logic, which was first developed by Prior in 1957 to facilitate precise reasoning about time for use in solving philosophical problems [Pri57]. Pnueli’s adaptation, which came to be known as Linear Temporal Logic (LTL), was tailored to have the necessary expressiveness to specify temporal dependencies in programs. Pnueli’s LTL was originally used over infinite traces. More recently LTL interpreted over finite traces (LTLf) was proposed in [DGV+13] and shown to be a useful specification language [vDAPS09, BFM06, Gab04, PBvdA10, Wil99].
@@
Regular expressions over propositional formulas (REf), where a propositional formula $\varphi$ serves as a representation of the set of propositional interpretations that satisfy $\varphi$, are a more expressive alternative to LTLf, with regular expressions being equivalent in expressiveness to monadic second-order logic and LTLf being equivalent in expressiveness to monadic first-order logic. Nevertheless, REf lacks direct constructs for negation and conjunction. These can be added by allowing complementation and intersection of regular expressions, but this results in nonemptiness checking being of nonelementary complexity even for star-free regular expressions.
@@
Linear Dynamic Logic on Finite Traces (LDLf) is a temporal logic that extends propositional logic with operators that use regular expressions to express temporal constraints [DGV+13]. It has the formal expressiveness of REf, while having constructs for negation and conjunction. Notably, satisfiability checking for LDLf formulas is PSPACE-complete, the same as for LTLf. Thus, LDLf is more expressive than LTLf at no additional cost [DGV+13]. Pure-Past Linear Dynamic Logic on Finite Traces (PLDLf) is a variant of LDLf that allows for the expression of present and past temporal constraints [DGDSF+20], as opposed to LDLf, which can express present and future temporal constraints.
4. Introducing ALDLf
@@
Automata Linear Dynamic Logic on Finite Traces (ALDLf) is a variant of LDLf that extends propositional logic with operators that use nondeterministic finite automata (NFA) to express temporal constraints. It also provides direct constructs for past modalities. The use of NFAs rather than regular expressions offers two advantages. The first advantage is that NFAs are exponentially more succinct than regular expressions [GH09, GH15]. Satisfiability checking of ALDLf formulas is PSPACE-complete, so this succinctness comes at no cost in complexity. The second advantage is that sometimes NFAs can be a more convenient form of expression than regular expressions. It is sometimes easier in practice to describe behavior using a state-based formalism, as can be observed by the the use of such formalism in industrial language. For example, SystemVerilog Assertions (SVA) uses local variables [VR05], which are a form of state.
@@
Moreover, ALDLf and its satisfiability-checking procedure can be used for LDLf, by converting an LDLf formula to an equivalent ALDLf formula. This can be done by converting each occurrence of regular expression in the LDLf formula to an equivalent NFA using a standard algorithm, e.g., Thompson’s construction [Tho68]. Having an alternative method to perform LDLf satisfiability checking to the one provided in [DGV+13] is advantageous because there are syntactially valid LDLf formulas for which the construction of the transition function for the AFW provided in [DGV+13] may not terminate. A simple example is $\varphi=\langle p^{**}\rangle\textbf{false}$, where $p$ is an atomic proposition. It is easy to see that following the AFW transition function construction provided in Theorem 15 of [DGV+13] on $\varphi$ results in a loop. Nevertheless, converting $\varphi$ to an ALDLf formula and performing the satisfiability checking procedure provided here is well founded.
@@ 4.2. ALDLf Semantics
If the conditions in 4.1 below hold, then an ALDLf formula is true within a trace $\pi\in(2^{AP})^{+}$ at position $i$, where $0\leq i<|\pi|$, (denoted $\pi,i\models\varphi$), and an ALDLf path automaton $(R,T,\Delta,r,G)$ is satisfied by the subsequence of $\pi$ from position $i$ to position $j$ (denoted $\pi,i,j\models(R,T,\Delta,r,G)$ if the conditions in 4.2 hold:
@@
Intuitively, a path automaton divides a trace into a prefix and a remainder (there may be more than one such division). In the case of a formula of the form $\langle U\rangle\varphi$, a trace $\pi$ will satisfy this formula just in case thatthere exists a prefix of $\pi$ that is in the language of $U$ such that the remainder of $\pi$ is in the language of $\varphi$. In the case of a formula of the form $[U]\varphi$, a trace $\pi$ will satisfy this formula just in case that for every prefix of $\pi$ that is in the language of $U$ the remainder of $\pi$ is in the language of $\varphi$. Note that $\langle U\rangle\varphi\equiv\neg[U]\neg\varphi$ and $[U]\varphi\equiv\neg\langle U\rangle\neg\varphi$.
A trace $\pi$ is in the language of an ALDLf formula $\varphi$, denoted $\pi\models\varphi$, if $\pi,0\models\varphi$. If $\zeta$ is a propositional ALDLf formula, the notation $\pi_{i}\models\zeta$ may be used as a synonym for $\pi,i\models\zeta$, in order to emphasize the fact that the truth value of $\zeta$ at $i$ depends only on the contents of $\pi_{i}$. An ALDLf formula $\varphi$ is in negation normal form if all instances of the negation operator ($\neg$) that occur in $\varphi$, including those that occur in tests that are included in a path automaton’s alphabet, are applied to atomic propositions. An ALDLf formula can be converted to negation normal form by replacing all instances of abbreviations with their definitions, converting all instances of $\neg[U]\varphi$ to $\langle U\rangle\neg\varphi$ and all instances of $\neg\langle U\rangle\varphi$ to $[U]\neg\varphi$, applying DeMorgan’s rules to make all negations appear as inwardly nested as possible, and then eliminating all double negations. If $\varphi$ is a formula, then nnf($\varphi$) is the result of performing this procedure on $\varphi$.
4.3. Fischer-Ladner Closure for ALDLf
As we have just seen, the satisfaction relation for ALDLf is defined recursively, i.e., the satisfaction of a formula is defined in terms of the satisfaction of other formulas with the exception of some base cases. This is similar to how satisfaction is defined for LTL. Our strategy to translate ALDLf formulas to automata will be to use states of the automaton to represent formulas that are “reachable” during the recursive evaluation of satisfaction, which is similar to how a similar translation is done for LTL. It is necessary to know in advance which formulas may need to be evaluated in order to construct the automaton. In the case of an LTL formula $\varphi$, the subformulas of $\varphi$ are sufficient. But in the case of modal logics such as ALDLf, the concept of subformula is too weak and so we need to use something more broad. The Fischer-Ladner closure, introduced in [FL79] for use in Propositional Dynamic Logic (PDL), is a generalization of subformulas for modal logics. Here we define the Fischer-Ladner closure for ALDLf.
@@ 4.4. Example
We can use these path automata to create a formula that defines the language of traces in which $a$ being true anywhere in the trace implies that $a$ is true at the first position in the trace: $\varphi=[\alpha_{3}]\langle\alpha_{2}\rangle[\alpha_{1}]\textbf{false}$. For the purpose of explanation, we break down $\varphi$ into these subformulas:
@@
5. Two-way Alternating Automata on Finite Words (2AFW)
In order to serve as an intermediate step in the translation of formulas of ALDLf to equivalent NFAs, we use a two-way alternating automaton on finite words (2AFW). A useful history and taxonomy of two-way alternating automata is provided by Kapoutsis and Zakzok [KZ21]. The 2AFW we define here is similar to that of Geffert and Okhotin [GO14]. The translation of ALDLf formula to NFA will first convert an ALDLf formula to an equivalent 2AFW and then convert the result to an equivalent NFA.